### v3

# An analogue of Ramanujan’s identity for Bernoulli-Carlitz numbers

Abstract.
In his second notebook, Ramanujan discovered the following identity for the special values of $\zeta(s)$ at the odd positive integers
$\alpha^{-m}$$\left\{\dfrac{1}{2}\,\zeta(2m+1)+\sum_{n=1}^{\infty}\dfrac{n^{-2m-1}}{e^{2\alpha n}-1}\right\}$
$\quad\quad\quad\quad\quad\quad\quad-(-\beta)^{-m}\,\left\{\dfrac{1}{2}\,\zeta(2m+1)+\sum_{n=1}^{\infty}\dfrac{n^{-2m-1}}{e^{2\beta n}-1}\right\}$
$=2^{2m}\sum_{k=0}^{m+1}\dfrac{\left(-1\right)^{k-1}B_{2k}\,B_{2m-2k+2}}{\left(2k\right)!\left(2m-2k+2\right)!}\,\alpha^{m-k+1}\beta^{k},$
where $\alpha$ and $\beta$ are positive numbers such that $\alpha\beta=\pi^{2}$ and $m$ is a positive integer. As shown by Berndt in the viewpoint of general transformation of analytic Eisenstein series, it is a natural companion of Euler’s famous formula for even zeta values.
In this note, we prove an analogue of the above Ramanujan’s identity in the functions fields setting, which involves the Bernoulli-Carlitz numbers.
Key words and phrases:
Ramanujan’s identity, Zeta value, Function fields, Goss zeta function, Bernoulli-Carlitz number
2010 Mathematics Subject Classification:
11R58, 11G09, 11M06, 11B68
1. Introduction
Let
$\zeta(s)=\sum_{n=1}^{\infty}\frac{1}{n^{s}},~{}~{}\textrm{Re}(s)>1$
be the Riemann zeta function. It can be analytically continued to the whole complex plane except for a single pole at $s=1$.
Around 1742, Euler obtained the special values of $\zeta(s)$ at the positive even integers
(1.1) $\zeta(2m)=(2\pi)^{2m}\frac{(-1)^{m}B_{2m}}{2(2m)!},$
where $B_{2m}$ are the Bernoulli numbers defined by the generating function
(1.2) $\frac{z}{e^{z}-1}=\sum_{m=0}^{\infty}B_{m}\frac{z^{m}}{m!}.$
Then Ramanujan discovered the following formula for the special values of $\zeta(s)$ at the odd positive integers (see Entry 21(i) in Ramanujan’s second notebook [3, p. 173] or his lost notebook [1, p. 319–320, formula (28)]).
Theorem 1.1 (Ramanujan’s formula for $\zeta(2m+1)$).
If $\alpha$ and $\beta$ are positive numbers such that $\alpha\beta=\pi^{2}$ and if $m$ is a positive integer, then we have
(1.3) $\alpha^{-m}$$\left\{\dfrac{1}{2}\,\zeta(2m+1)+\sum_{n=1}^{\infty}\dfrac{n^{-2m-1}}{e^{2\alpha n}-1}\right\}$
$\quad\quad\quad\quad\quad\quad\quad-(-\beta)^{-m}\,\left\{\dfrac{1}{2}\,\zeta(2m+1)+\sum_{n=1}^{\infty}\dfrac{n^{-2m-1}}{e^{2\beta n}-1}\right\}$
$=2^{2m}\sum_{k=0}^{m+1}\dfrac{\left(-1\right)^{k-1}B_{2k}\,B_{2m-2k+2}}{\left(2k\right)!\left(2m-2k+2\right)!}\,\alpha^{m-k+1}\beta^{k}.$
In 1925, Marulkar [16] gave the first proof of (1.3), but he did not aware that this formula can be found in Ramanujan’s Notebooks. Then (1.3) has been rediscovered and studied by Grosswald in [8, 9]. In 1977, Berndt [2, Theorem 2.2] showed that both Euler’s formula (1.1) and Ramanujan’s formula (1.3) are special cases of a general transformation formula of analytic Eisenstein series, thus they are natural companions of each other. And in 2011, Gun, Murty and Rath [14] gave a modern interpretation of (1.3), that is, it encodes fundamental transformation properties of Eisenstein series on the full modular group and their Eichler integrals. This approach has been further extended by Berndt and Straub [4, Section 5] to weight $12k+$ 1 Eisenstein series of level 2 through secant Dirichlet series. Recently, Chavan [7] presented an elementary proof of (1.3), which relies on a Mittag-Leffler type expansion of hyperbolic contangent function
(1.4) $\coth\left(\pi x\right)=\dfrac{1}{\pi x}+\dfrac{2x}{\pi}\sum_{k=1}^{\infty}\dfrac{1}{x^{2}+k^{2}}$
for $x\in\mathbb{C}\setminus{\{0\}}$ and Euler’s identity for even zeta values (1.1).
In this note, we shall investigate the above Ramanujan’s identity in the functions fields setting. In this situation, an analogy of the Riemann zeta function $\zeta(s)$ was introduced by Goss [12] in 1979. Let $q=p^{k}$ be a power of a prime number $p$, $\mathbb{F}_{q}$ be the finite field with $q$ elements and $\mathbb{F}_{q}^{*}$ be its multiplicative group. Let $R=\mathbb{F}_{q}[T]$ be the polynomial ring with one variable over the finite field $\mathbb{F}_{q}$ and $K=\mathbb{F}_{q}(T)$ be the rational function field. Let $K_{\infty}=\mathbb{F}_{q}((\frac{1}{T}))$ be the completion of $K$ at the infinite place $\infty=\left(\frac{1}{T}\right)$, $K_{\infty}^{*}=K_{\infty}\setminus\{0\}$ be the multiplicative group of the field $K_{\infty}$ and $\mathbb{C}_{\infty}$ be the completion of an algebraic closure of $K_{\infty}$. Let $R_{1}$ be the set of monic polynomials in $R,$ which is an analogue of the set of positive integers. Goss considered the infinite series
(1.5) $\zeta_{\infty}(s)=\sum_{A\in R_{1}}A^{-s}$
for $s\in\mathbb{S}=K_{\infty}^{*}\times\mathbb{Z}_{p},$ where $\mathbb{Z}_{p}$ is the ring of $p$-adic integers, and he also showed that it is an analytic function on $\mathbb{S}$. For details, we refer to Goss’ original article [12].
Let
$[j]=T^{q^{j}}-T,\quad D_{j}=[j]{[j-1]}^{q}\cdots{[1]}^{q^{j-1}}$
with $D_{0}=1$. The Carlitz exponential is defined by
(1.6) $e(z)=\sum_{j=0}^{\infty}\frac{z^{q^{j}}}{D_{j}}.$
It is an entire $\mathbb{F}_{q}$-linear function and has the following infinite product expansion
(1.7) $e(z)=e_{\tilde{\pi}R}(z)=z\prod_{0\neq\lambda\in\tilde{\pi}R}\left(1-\frac{z}{\lambda}\right),$
where $\tilde{\pi}\in\mathbb{C}_{\infty}$ is an analogy of the classical $2\pi i~{}(i=\sqrt{-1})$, as it is a fundamental period of Carlitz’s exponential $e(z)$. In 1941, Wade [19] proved that $\tilde{\pi}$ is transcendental over $K$.
Let $\Omega=\mathbb{C}_{\infty}\setminus K_{\infty}$ be Drinfeld’s upper half-plane, which is an analogue of the classical upper half-plane $H=\{\tau\in\mathbb{C}~{}|~{}\textrm{Im}(\tau)>0\}.$ Recently, by applying $e(z)$, Hamahata [15] introduced a function fields analogue of Carlitz’s generalized Dedekind sums,
(1.8) $t_{r}^{(n)}(a,c)=\sum_{0\neq\mu\in R/cR}e(\tilde{\pi}\mu/c)^{-n-1+r}e(\tilde{\pi}a\mu/c)^{-r}$
for $a,c\in R\setminus\{0\}$ and $r\in\{0,1,\ldots,n+1\}$. It enables him to establish the Apostol–type transformation formulas for the following Lambert series in function fields,
(1.9) $\xi_{n}(z)=\sum_{a\in R_{1}}\frac{1}{a^{n}}\frac{1}{e(\tilde{\pi}az)},~{}~{}z\in\Omega.$
(See [15, Section 4]).
For a nonnegative integer $m$ with $q$-ary expansion $m=\sum_{j=0}^{w}c_{j}q^{j}~{}(0\leq c_{j}<q),$ the Carlitz factorial $\Gamma_{m}$ is
$\Gamma_{m}=\prod_{j=0}^{w}D_{j}^{c_{j}}.$
Then in analogy with the definition of the classical Bernoulli numbers (1.2), the Bernoulli-Carlitz numbers $BC_{m}\in K~{}(n=0,1,2,\ldots)$ are defined by the generating function
(1.10) $\frac{z}{e(z)}=\sum_{m=0}^{\infty}\frac{BC_{m}}{\Gamma_{m}}{z^{m}}.$
As in the classical case, the Bernoulli-Carlitz numbers have many deep applications in the arithmetic of function fields, such as Kummer’s and Herbrand’s criterions for the class groups of cyclotomic function fields (see [13, Section 9.2], [18, Section 5.2] or [10, 11]). In 1930s, Carlitz [5, 6] showed an analogue of Euler’s formula (1.1), that is, the special values of $\zeta_{\infty}(s)$ at the “even” positive integers $m=(q-1)j~{}(j\in\mathbb{N})$ are given by
(1.11) $\zeta_{\infty}(m)=\tilde{\pi}^{m}\frac{BC_{m}}{\Gamma_{m}}.$
For details on the special values of $\zeta_{\infty}(s)$, we refer to Thakur’s work [17].
Now suppose $q$ is a power of an odd prime and by understanding that $BC_{0}=\Gamma_{0}=1$. In this note, we shall prove the following function fields analogue of Ramanujan’s identity (1.3).
Theorem 1.2.
If $\alpha$ and $\beta$ are elements in $\mathbb{C}_{\infty}$ such that $\alpha\beta=\tilde{\pi}^{2}$ and if $j$ is a positive integer, then we have
$\alpha^{-(\frac{q-1}{2})(j-1)-(q-2)}$ $\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\alpha A)}+\beta^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\beta A)}$
(1.12) $=\sum_{k=0}^{j+1}\frac{BC_{(q-1)k}}{\Gamma_{(q-1)k}}\frac{BC_{(q-1)(j-k+1)}}{\Gamma_{(q-1)(j-k+1)}}\alpha^{(\frac{q-1}{2})k}\beta^{(\frac{q-1}{2})(j-k+1)}.$
Remark 1.3.
By comparing (1.12) with (1.3), we see that the terms
$\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\alpha A)}\quad\text{and}\quad\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\beta A)}$
are analogies of
$\sum_{n=1}^{\infty}\dfrac{n^{-2m-1}}{e^{2\alpha n}-1}\quad\text{and}\quad\sum_{n=1}^{\infty}\dfrac{n^{-2m-1}}{e^{2\beta n}-1},$
respectively.
2. Proof of Theorem 1.2
Our proof is inspired by Chavan [7] for his elementary proof of Ramanujan’s formula (1.3).
By (1.7), we get
(2.1) $e(z)$ $=e_{\tilde{\pi}R}(z)$
$=z\prod_{0\neq\lambda\in\tilde{\pi}R}\left(1-\frac{z}{\lambda}\right)$
$=z\prod_{0\neq A\in R}\left(1-\frac{z}{\tilde{\pi}A}\right)$
$=z\prod_{A\in R_{1}}\left(1-\left(\frac{z}{\tilde{\pi}A}\right)^{q-1}\right),$
the last identity follows from the polynomial factorization
$x^{q-1}-y^{q-1}=\prod_{\xi\in\mathbb{F}_{q}^{*}}(x-\xi y).$
From (1.6),
$e^{\prime}(z)=1.$
So
(2.2) $\frac{1}{e(z)}$ $=\frac{e^{\prime}(z)}{e(z)}$
$=\frac{1}{z}+\sum_{A\in R_{1}}\frac{-\frac{(q-1)z^{q-2}}{(\tilde{\pi}A)^{q-1}}}{1-\left(\frac{z}{\tilde{\pi}A}\right)^{q-1}}$
$=\frac{1}{z}-z^{q-2}\sum_{A\in R_{1}}\frac{1}{z^{q-1}-(\tilde{\pi}A)^{q-1}}.$
Thus for any $\alpha\in\Omega$ and $B\in R$, we have
(2.3) $\frac{1}{e(\alpha B)}=\frac{1}{\alpha B}-(\alpha B)^{q-2}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}.$
Note that, by (2.3) and (1.5), for any positive integer $j,$ we have
(2.4) $\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
$=\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}}\frac{1}{(\alpha B)^{q-2}}\left(\frac{1}{\alpha B}-\frac{1}{e(\alpha B)}\right)$
$=\alpha^{(\frac{q-1}{2})(j+1)}\left(\frac{1}{\alpha^{(q-1)(j+1)}}\zeta_{\infty}\left((q-1)(j+1)\right)\right.$
$\left.\qquad\qquad\qquad\qquad\qquad\quad-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j+(q-2)}e(\alpha B)}\right).$
Then by applying (1.11) and notice that $\alpha\beta=\tilde{\pi}^{2}$, we further get
(2.5) $\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
$=\alpha^{(\frac{q-1}{2})(j+1)}\frac{1}{\alpha^{(q-1)(j+1)}}\tilde{\pi}^{(q-1)(j+1)}\frac{BC_{(q-1)(j+1)}}{\Gamma_{(q-1)(j+1)}}$
$\quad-\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j+(q-2)}e(\alpha B)}$
$=\beta^{(\frac{q-1}{2})(j+1)}\frac{BC_{(q-1)(j+1)}}{\Gamma_{(q-1)(j+1)}}-\alpha^{-(\frac{q-1}{2})(j-1)-(q-2)}$
×$\quad\times\sum_{B\in R_{1}}\frac{B^{-(q-1)j-(q-2)}}{e(\alpha B)}.$
On the other hand, a direct calculation shows that
(2.6) $\quad\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
$=-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{q-1}}$
$\quad+\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j-1)}(\tilde{\pi}A)^{q-1}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
$=-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{q-1}}$
$\quad-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j-1)}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)2}}$
$\quad+\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j-2)}(\tilde{\pi}A)^{(q-1)2}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
=$\cdots$= ⋯
[…]
$=-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)j}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{q-1}}$
$\quad-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j-1)}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)2}}$
[…]
$\quad-\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{q-1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}$
$\quad+\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
$=-\sum_{k=1}^{j}\left(\sum_{B\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j-k+1)}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)k}}\right)$
$\quad+\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
$=-\sum_{k=1}^{j}\frac{\zeta_{\infty}\left((q-1)(j-k+1)\right)}{\alpha^{(q-1)(j-k+1)}}\frac{\zeta_{\infty}\left((q-1)k\right)}{\tilde{\pi}^{(q-1)k}}$
$\quad+\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}},$
the last identity follows from (1.5).
Then multiplying both sides of the above equation by $\alpha^{(\frac{q-1}{2})(j+1)},$ notice (1.11) and the relation $\alpha\beta=\tilde{\pi}^{2}$, we have
(2.7) $\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\alpha B)^{(q-1)(j+1)}\left((\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}\right)}$
$=-\sum_{k=1}^{j}\frac{\zeta_{\infty}\left((q-1)(j-k+1)\right)}{\tilde{\pi}^{(q-1)(j-k+1)}}\frac{\zeta_{\infty}\left((q-1)k\right)}{\tilde{\pi}^{(q-1)k}}\alpha^{(\frac{q-1}{2})k}\beta^{(\frac{q-1}{2})(j-k+1)}$
$\quad+\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
$=-\sum_{k=1}^{j}\frac{BC_{(q-1)(j-k+1)}}{\Gamma_{(q-1)(j-k+1)}}\frac{BC_{(q-1)k}}{\Gamma_{(q-1)k}}\alpha^{(\frac{q-1}{2})k}\beta^{(\frac{q-1}{2})(j-k+1)}$
$\quad+\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}.$
By comparing with (2.5) we get
(2.8) $\quad-\alpha^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{B\in R_{1}}\frac{B^{-(q-1)j-(q-2)}}{e(\alpha B)}$
$=-\sum_{k=0}^{j}\frac{BC_{(q-1)(j-k+1)}}{\Gamma_{(q-1)(j-k+1)}}\frac{BC_{(q-1)k}}{\Gamma_{(q-1)k}}\alpha^{(\frac{q-1}{2})k}\beta^{(\frac{q-1}{2})(j-k+1)}$
$\quad+\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}.$
To compute the last term, note that $\alpha\beta=\tilde{\pi}^{2}$, we have
(2.9) $\quad\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
$=\frac{1}{\tilde{\pi}^{(q-1)j}}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{A^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-\alpha^{\frac{q-1}{2}}\beta^{\frac{q-1}{2}}A^{q-1}}$
$=\frac{1}{\tilde{\pi}^{(q-1)j}}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{\alpha^{\frac{q-1}{2}}A^{(q-1)j}}\frac{1}{\alpha^{\frac{q-1}{2}}B^{q-1}-\beta^{\frac{q-1}{2}}A^{q-1}}$
$=\frac{1}{\alpha^{(\frac{q-1}{2})j}\beta^{(\frac{q-1}{2})j}}\sum_{A\in R_{1}}\frac{\beta^{\frac{q-1}{2}}}{\alpha^{\frac{q-1}{2}}}\frac{1}{A^{(q-1)j}}\frac{1}{\tilde{\pi}^{q-1}B^{q-1}-\beta^{q-1}A^{q-1}}$
$=-\frac{\beta^{(\frac{q-1}{2})j}}{\alpha^{(\frac{q-1}{2})j}}\frac{\beta^{\frac{q-1}{2}}}{\alpha^{\frac{q-1}{2}}}\sum_{A\in R_{1}}\frac{1}{(\beta A)^{(q-1)j}}\frac{1}{(\beta A)^{q-1}-({\tilde{\pi}B})^{q-1}}.$
Then multiplying both sides of the above equation by $\alpha^{(\frac{q-1}{2})(j+1)}$, notice (2.3) and (1.5), we get
(2.10) $\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
$=-\beta^{(\frac{q-1}{2})(j+1)}\sum_{A\in R_{1}}\frac{1}{(\beta A)^{(q-1)j}}\frac{1}{(\beta A)^{q-2}}\left(\frac{1}{\beta A}-\frac{1}{e(\beta A)}\right)$
$=-\beta^{(\frac{q-1}{2})(j+1)}\sum_{A\in R_{1}}\frac{1}{(\beta A)^{(q-1)j}}\frac{1}{(\beta A)^{q-2}}\left(\frac{1}{\beta A}-\frac{1}{e(\beta A)}\right)$
$=-\beta^{(\frac{q-1}{2})(j+1)}\left(\frac{1}{\beta^{(q-1)(j+1)}}\zeta_{\infty}\left((q-1)(j+1)\right)\right.$
$\qquad\qquad\qquad\qquad\qquad\quad-\left.\sum_{A\in R_{1}}\frac{1}{(\beta A)^{(q-1)j+(q-2)}e(\beta A)}\right).$
Similarly with the reasoning of (2.5), by applying (1.11) to the above equation and notice that $\alpha\beta=\tilde{\pi}^{2}$, we see that
(2.11) $\quad\alpha^{(\frac{q-1}{2})(j+1)}\sum_{B\in R_{1}}\sum_{A\in R_{1}}\frac{1}{(\tilde{\pi}A)^{(q-1)j}}\frac{1}{(\alpha B)^{q-1}-(\tilde{\pi}A)^{q-1}}$
$=-\beta^{(\frac{q-1}{2})(j+1)}\frac{1}{\beta^{(q-1)(j+1)}}\tilde{\pi}^{(q-1)(j+1)}\frac{BC_{(q-1)(j+1)}}{\Gamma_{(q-1)(j+1)}}$
$\quad+\beta^{(\frac{q-1}{2})(j+1)}\sum_{A\in R_{1}}\frac{1}{(\beta A)^{(q-1)j+(q-2)}e(\beta A)}$
$=-\alpha^{(\frac{q-1}{2})(j+1)}\frac{BC_{(q-1)(j+1)}}{\Gamma_{(q-1)(j+1)}}$
$\quad+\beta^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\beta A)}.$
Finally, by substituting the above equation into (2.8) we obtain
(2.12) $\alpha^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{B\in R_{1}}\frac{B^{-(q-1)j-(q-2)}}{e(\alpha B)}$
$\quad+\beta^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\beta A)}$
$=\sum_{k=0}^{j+1}\frac{BC_{(q-1)(j-k+1)}}{\Gamma_{(q-1)(j-k+1)}}\frac{BC_{(q-1)k}}{\Gamma_{(q-1)k}}\alpha^{(\frac{q-1}{2})k}\beta^{(\frac{q-1}{2})(j-k+1)},$
which is the desired result.